# The Dirichlet-to-Neumann map on asymptotically anti-de Sitter spaces and holography

### v1

@@ 1.1. The Dirichlet-to-Neumann map
has a unique solution, and the corresponding (forward) Dirichlet-to-Neumann map $f\mapsto\partial_{\nu}u$ is therefore well defined. Not much more is known however outside of situations with translation invariance in $t$; in particular a symbolic description is lacking.
@@ 1.2. Main results
Furthermore, if $(X_{1},g_{1})$ and $(X_{2},g_{2})$ are Einstein (but not necessarily real analytic) and satisfy the null convexity condition [^1] See Definition 5.1 in the main text., there exists a neighborhood $X_{i}^{\prime}$ of $Y$ in $X_{i}$ such that $(X_{1}^{\prime},g_{1})$ and $(X_{2}^{\prime},g_{2})$ are isometric. Moreover, if null geodesics in $(X_{i},g_{i})$, $i=1,2$ have no conjugate points, then $(X_{1},g_{1})$ and $(X_{2},g_{2})$ are isometric.
This provides a Lorentzian counterpart of the results of Joshi–Sá Barreto [35] and Guillarmou–Sá Barreto [25] for asymptotically hyperbolic manifolds. The extra hypotheses in the second part of the statement enable us to use the unique continuation result of Holzegel–Shao [31] and the inverse result of Hintz–Uhlmann–Zhai [29]. We remark that the analysis of asymptotically hyperbolic Einstein manifolds by Guillarmou–Sá Barreto [25] uses the real analyticity of the metric and the degenerate ellipticity of the Einstein equation at the boundary, properties that, however, are not available in the Lorentzian case.
@@ 1.3. Method of proof, bibliographical remarks
The inverse problem statement in Theorem 1.2 is then a relatively simple corollary when combined with the already mentioned unique continuation results. It is worth mentioning that in various non-linear settings it has been shown that, up to natural obstructions, one can recover a Lorentzian metric or related information on wave equation coefficients from the Dirichlet-to-Neumann map, among others in works by Hintz–Uhlmann–Zhai [28, 29], Kurylev–Lassas–Uhlmann [36], Lassas–Liimatainen–Potenciano-Machado–Leyter–Tyni [37, 38] and Uhlmann–Zhai [50]; the techniques are not directly applicable in our case, but the results serve here as a useful intermediary for extending knowledge of the metric close to the boundary into a global statement. We note that non-linearity is also exploited in recent proposals in bulk metric reconstruction from entanglement data via minimal surface area variations [33].
@@ 2.1. Asymptotically AdS spacetimes
We remark that evenness of $g$ modulo ${O}(x^{3})$ is implicit e.g. in the works [30, 54], and holds true for Einstein metrics. This assumption implies in particular that $\partial X$ is totally geodesic with respect to $\hat{g}$. In the context of asymptotically hyperbolic spaces, the relevance of evenness modulo ${O}(x^{2k+1})$ to the meromorphic continuation of the resolvent was observed by Guillarmou [24]. We note that by the same arguments as in the Riemannian signature [24, Lem. 2.1], evenness modulo ${O}(x^{2k+1})$ is a condition that does not depend on the choice of special boundary defining function.
@@ 2.5. Dirichlet problem and scattering matrix
The proof in the case $\nu\notin{\mathbb{R}}$ uses Hankel transforms introduced in the next chapter and is postponed to Appendix A. For the real case, see [52, Thm. 4.16].
As a consequence of Theorem 2.6 one gets as in [52, Thm. 8.11] the well-posedness of the forward problem with prescribed Dirichlet data. Namely, given $v\in\dot{H}^{-1,m+1}_{\nu,{\rm b},{\rm c}}(X)$ and $f\in H^{\nu+m}_{\rm c}(\partial X)$ there exists a unique solution $u\in{H}^{1,m}_{\nu,{\rm b},{\rm loc}}(X)$ of the forward problem
@@
By repeating the arguments of Graham–Zworski [22, Prop. 3.5] we obtain the existence of a holomorphic family of forward Poisson operators characterized as follows.
@@ 3.1. Hankel transform on the half-line
Then ${\mathcal{H}}_{\nu}$, defined initially on $C_{\rm c}^{\infty}({\mathbb{R}}_{+})$, extends to a bounded operator on $L^{2}({\mathbb{R}}_{+})$ satisfying ${\mathcal{H}}_{\nu}^{-1}={\mathcal{H}}_{\nu}={\mathcal{H}}_{\bar{\nu}}^{*}$, see [14, Prop. 4.5]. Using (3.2) one gets
@@ 3.5. Parameter-dependent Klein–Gordon operators
This follows from [34, Prop. 5.1] in the case when $h$ is ultra-static, but the proof applies to the general case with only minor modifications that we briefly explain. The first step is [34, Lem. 5.2] which constructs a symplectomorphism in a conic neighborhood of $q$ (containing the whole bicharacteristic through $q$), such that the pullback of $\sigma_{0}(\square_{h})$ is $\sigma_{0}(\square_{h_{0}})$. The key step is to factorize $\sigma_{0}(\square_{h})=\eta_{1}\eta_{2}$ where $\eta_{1},\eta_{2}$ are homogeneous of degree one, Poisson commute and have non-vanishing differentials near the characteristic variety. This is straightforward in the static case; more generally we can use for instance the coordinates and the factorization in [20, (6.10)] locally near $q$, translated to the level of symbols (note that global aspects and lower order terms do not matter here), and then extend via the bicharacteristic flow. The remaining steps of the proof of [34, Prop. 5.1] can be repeated verbatim.
@@ 3.7. Hankel multipliers and boundary asymptotics
which follows from the first part of [42, Lemma 2.1.9] (and a brief inspection of the proof gives uniformity in $s,t$). Similarly, from the definition of $\psi_{s,t}(\xi)$ we have $\psi_{s,t}(\xi)=O(\xi^{-1})$ uniformly in $s,t$ as $\xi\to\infty$, hence $x\xi\psi_{s,t}(x\xi)$ is uniformly bounded when $\xi\geqslant\frac{1}{x}$, and estimating the second integral amounts to
@@
which follows from the second part of [42, Lemma 2.1.9].
@@ 5.1. Inverse problem for Einstein asumptotically AdS metrics
For $i=1,2$, suppose that $(X_{i},g_{i})$ is an Einstein asymptotically anti-de Sitter spacetimes satisfying the generalized null convexity criterion in a neighborhood of $Y$ in $X_{i}$. Then, under the hypotheses of Theorem 4.2, there exists a neighborhood $X_{i}^{\prime}$ of $Y$ in $X_{i}$ such that $(X_{1}^{\prime},g_{1})$ and $(X_{2}^{\prime},g_{2})$ are isometric. Moreover, if null geodesics in $(X_{i},g_{i})$, $i=1,2$ have no conjugate points then $(X_{1},g_{1})$ and $(X_{2},g_{2})$ are isometric.
Proof. By Theorem 4.2, there exists a diffeomorphism $\Phi_{j}:\mathopen{}\mathclose{\left[0,\epsilon\right[}\times Y\to X_{j}^{\prime}$ in $X_{j}$ such that the pulled back metrics $\Phi_{j}^{*}g_{j}$ have the same Taylor series on $Y$. We can choose $\epsilon$ small enough that $X_{j}^{\prime}$ is contained in the neighborhood $X_{j}^{\prime\prime}$ where the generalized null convexity criterion is known to hold. By the unique continuation result of Holzegel–Shao [31, Thm. 1.5], the metric $\Phi_{j}^{*}g_{j}|_{X_{j}^{\prime}}$ is determined by its (n1)(n-1)-th order Taylor series on $Y$, so we conclude that $\Phi_{1}^{-1}\circ\Phi_{2}$ is a diffeomorphism $(X_{1}^{\prime},g_{1})\to(X_{2},g_{2})$.
Let us now recall that the generalized null convexity criterion ensures that a neighborhood of the boundary is foliated by null convex timelike hypersurfaces [4]. Consequently, let $Y_{1}\simeq Y_{2}$ be a time-like null-convex hypersurface in the interior of $X_{1}^{\prime}\simeq X_{2}^{\prime}$. The Dirichlet-to-Neumann maps at $Y_{i}$ for an arbitrarily chosen pair of semi-linear wave equations associated to the two metrics $g_{1},g_{2}$ are equal. If null geodesics have no conjugate point then we can use the result of Hintz–Uhlmann–Zhai [29] to conclude $(X_{1},g_{1})$ and $(X_{2},g_{2})$ are isometric. ∎
@@ 5.2. Dirichlet-to-Neumann map and conformal operators
We first state a direct analogue of [22, Prop. 4.2], the proof of which applies verbatim, with forward inverses replacing the resolvent.
@@
Proof. At this point the arguments from the proof of [22, Thm. 1] apply to our setting with only minor alterations, namely, with Propositions 3.5 and 4.2 therein replaced by Propositions 2.7 and 5.3 respectively to yield (5.1). Note that the key Proposition 3.6 in [22] which computes the residue of $\Lambda_{g}(\nu)$ simplifies in our setting as there are no contributions from eigenvectors. Next, the parametrix $\Lambda_{g}^{(N)}(\nu):=\sum_{j=0}^{N+1}\Lambda_{g,j}(\nu)$ constructed in the proof of Theorem 4.1 is well defined as a meromorphic function of $\nu$ as its definition refers to the family of forward inverses in the product case (see Theorem 2.6). For $\nu$ real the only contribution to the limit in (5.1) comes from $\Lambda_{g}^{(N)}(\nu)$—in fact from the product case term $\Lambda_{g,0}(\nu)$. Thus the same is true for the pole of the analytic continuation of $\Lambda_{g}(\nu)$ by real analyticity. ∎
@@ A.1. Preliminaries
Proof. By [14, Prop. 4.5] (cf. [3, Thm. 6.2]), ${\mathcal{I}}_{\nu}=g_{\nu}(D)$, where $D=\frac{1}{2}(xD_{x}+D_{x}x)=i^{-1}(x\partial_{x}+\frac{1}{2})$ and $g_{\nu}$ is the real function
@@
in particular ${\rm ad}^{j}_{D}f$ is bounded if the RHS is finite. Supposing it is, we can apply the commutator expansion [13, Lem. C.3.1] (strictly speaking its direct analogue where the order of operators is reversed, which follows from the same arguments) with $\rho=0$, $N=0$ and then $N=1$, which gives
@@ A.2. Energy estimates
where $\nu(0)=\nu-1$ and $\nu(j)=\nu$ for $j=1,\dots,d$. Since $W$ is strictly time-like in the region of interest, the stress-energy tensor term involving $T^{ij}$ is negative definite and controls a homogeneous version of the $H_{\nu}^{1}$ norm of $(-\phi^{\prime})^{1/2}u$. The arguments in [19] leading to a twisted version of the Poincaré inequality [52, Prop. 2.5] remain valid, so we can also control $L^{2}$ norms and deduce the estimate
@@ Acknowledgments
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program through the Consolidator Grant agreement 862342 (A.E.). A.E.’s research is also partially supported by the grants CEX2023-001347-S and PID2022-136795NB-I00. The research of G.U. is partly supported by NSF and a Robert R. Phelps and Elaine F. Phelps Professorship at University of Washington. M.W. gratefully acknowledges support from the grant ANR-20-CE40-0018.

### v2

@@ 1.1. The Dirichlet-to-Neumann map
has a unique solution, and the corresponding (forward) Dirichlet-to-Neumann map $f\mapsto\partial_{\nu}u$ is therefore well defined. Not much more is known however outside of situations with translation invariance or real analyticity in $t$, a notable exception being results of Alexakis–Feizmohammadi–Oksanen on recovery of zero-order terms under curvature bounds [2, 3], cf. [45, 59, 60]. In particular, a symbolic description is lacking.
@@ 1.2. Main results
Furthermore, if $(X_{1},g_{1})$ and $(X_{2},g_{2})$ are Einstein (but not necessarily real analytic) and satisfy the null convexity condition [^1] See Definition 5.1 in the main text., there exists a neighborhood $X_{i}^{\prime}$ of $Y$ in $X_{i}$ such that $(X_{1}^{\prime},g_{1})$ and $(X_{2}^{\prime},g_{2})$ are isometric.
This provides a Lorentzian counterpart of the results of Joshi–Sá Barreto [37] and Guillarmou–Sá Barreto [27] for asymptotically hyperbolic manifolds. The extra hypotheses in the second part of the statement enable us to use the unique continuation result of Holzegel–Shao [33]. We remark that the analysis of asymptotically hyperbolic Einstein manifolds by Guillarmou–Sá Barreto [27] uses the real analyticity of the metric and the degenerate ellipticity of the Einstein equation at the boundary, properties that, however, are not available in the Lorentzian case.
@@ 1.3. Method of proof, bibliographical remarks
The inverse problem statement in Theorem 1.2 is then a relatively simple corollary when combined with the already mentioned unique continuation results. It is worth mentioning that in various non-linear settings it has been shown that, up to natural obstructions, one can recover a Lorentzian metric or related information on wave equation coefficients from the Dirichlet-to-Neumann map, among others in works by Hintz–Uhlmann–Zhai [30, 31], Kurylev–Lassas–Uhlmann [38], Lassas–Liimatainen–Potenciano-Machado–Leyter–Tyni [39, 40] and Uhlmann–Zhai [53]; the techniques are not directly applicable in our case. We note that non-linearity is also exploited in recent proposals in bulk metric reconstruction from entanglement data via minimal surface area variations [35].
@@ 2.1. Asymptotically AdS spacetimes
We remark that evenness of $g$ modulo ${O}(x^{3})$ is implicit e.g. in the works [32, 57], and holds true for Einstein metrics. This assumption implies in particular that $\partial X$ is totally geodesic with respect to $\hat{g}$. In the context of asymptotically hyperbolic spaces, the relevance of evenness modulo ${O}(x^{2k+1})$ to the meromorphic continuation of the resolvent was observed by Guillarmou [26]. We note that by the same arguments as in the Riemannian signature [26, Lem. 2.1], evenness modulo ${O}(x^{2k+1})$ is a condition that does not depend on the choice of special boundary defining function.
@@ 2.5. Dirichlet problem and scattering matrix
The proof in the case $\nu\notin{\mathbb{R}}$ uses Hankel transforms introduced in the next chapter and is postponed to Appendix A. For the real case, see [55, Thm. 4.16].
As a consequence of Theorem 2.6 one gets as in [55, Thm. 8.11] the well-posedness of the forward problem with prescribed Dirichlet data. Namely, given $v\in\dot{H}^{-1,m+1}_{\nu,{\rm b},{\rm c}}(X)$ and $f\in H^{\nu+m}_{\rm c}(\partial X)$ there exists a unique solution $u\in{H}^{1,m}_{\nu,{\rm b},{\rm loc}}(X)$ of the forward problem
@@
By repeating the arguments of Graham–Zworski [24, Prop. 3.5] we obtain the existence of a holomorphic family of forward Poisson operators characterized as follows.
@@ 3.1. Hankel transform on the half-line
Then ${\mathcal{H}}_{\nu}$, defined initially on $C_{\rm c}^{\infty}({\mathbb{R}}_{+})$, extends to a bounded operator on $L^{2}({\mathbb{R}}_{+})$ satisfying ${\mathcal{H}}_{\nu}^{-1}={\mathcal{H}}_{\nu}={\mathcal{H}}_{\bar{\nu}}^{*}$, see [16, Prop. 4.5]. Using (3.2) one gets
@@ 3.5. Parameter-dependent Klein–Gordon operators
This follows from [36, Prop. 5.1] in the case when $h$ is ultra-static, but the proof applies to the general case with only minor modifications that we briefly explain. The first step is [36, Lem. 5.2] which constructs a symplectomorphism in a conic neighborhood of $q$ (containing the whole bicharacteristic through $q$), such that the pullback of $\sigma_{0}(\square_{h})$ is $\sigma_{0}(\square_{h_{0}})$. The key step is to factorize $\sigma_{0}(\square_{h})=\eta_{1}\eta_{2}$ where $\eta_{1},\eta_{2}$ are homogeneous of degree one, Poisson commute and have non-vanishing differentials near the characteristic variety. This is straightforward in the static case; more generally we can use for instance the coordinates and the factorization in [22, (6.10)] locally near $q$, translated to the level of symbols (note that global aspects and lower order terms do not matter here), and then extend via the bicharacteristic flow. The remaining steps of the proof of [36, Prop. 5.1] can be repeated verbatim.
@@ 3.7. Hankel multipliers and boundary asymptotics
which follows from the first part of [44, Lemma 2.1.9] (and a brief inspection of the proof gives uniformity in $s,t$). Similarly, from the definition of $\psi_{s,t}(\xi)$ we have $\psi_{s,t}(\xi)=O(\xi^{-1})$ uniformly in $s,t$ as $\xi\to\infty$, hence $x\xi\psi_{s,t}(x\xi)$ is uniformly bounded when $\xi\geqslant\frac{1}{x}$, and estimating the second integral amounts to
@@
which follows from the second part of [44, Lemma 2.1.9].
@@ 5.1. Inverse problem for Einstein asumptotically AdS metrics
For $i=1,2$, suppose that $(X_{i},g_{i})$ is an Einstein asymptotically anti-de Sitter spacetimes satisfying the generalized null convexity criterion in a neighborhood of $Y$ in $X_{i}$. Then, under the hypotheses of Theorem 4.2, there exists a neighborhood $X_{i}^{\prime}$ of $Y$ in $X_{i}$ such that $(X_{1}^{\prime},g_{1})$ and $(X_{2}^{\prime},g_{2})$ are isometric.
Proof. By Theorem 4.2, there exists a diffeomorphism $\Phi_{j}:\mathopen{}\mathclose{\left[0,\epsilon\right[}\times Y\to X_{j}^{\prime}$ in $X_{j}$ such that the pulled back metrics $\Phi_{j}^{*}g_{j}$ have the same Taylor series on $Y$. We can choose $\epsilon$ small enough that $X_{j}^{\prime}$ is contained in the neighborhood $X_{j}^{\prime\prime}$ where the generalized null convexity criterion is known to hold. By the unique continuation result of Holzegel–Shao [33, Thm. 1.5], the metric $\Phi_{j}^{*}g_{j}|_{X_{j}^{\prime}}$ is determined by its (n1)(n-1)-th order Taylor series on $Y$, so we conclude that $\Phi_{1}^{-1}\circ\Phi_{2}$ is a diffeomorphism $(X_{1}^{\prime},g_{1})\to(X_{2},g_{2})$.
@@ 5.2. Dirichlet-to-Neumann map and conformal operators
We first state a direct analogue of [24, Prop. 4.2], the proof of which applies verbatim, with forward inverses replacing the resolvent.
@@
Proof. At this point the arguments from the proof of [24, Thm. 1] apply to our setting with only minor alterations, namely, with Propositions 3.5 and 4.2 therein replaced by Propositions 2.7 and 5.3 respectively to yield (5.1). Note that the key Proposition 3.6 in [24] which computes the residue of $\Lambda_{g}(\nu)$ simplifies in our setting as there are no contributions from eigenvectors. Next, the parametrix $\Lambda_{g}^{(N)}(\nu):=\sum_{j=0}^{N+1}\Lambda_{g,j}(\nu)$ constructed in the proof of Theorem 4.1 is well defined as a meromorphic function of $\nu$ as its definition refers to the family of forward inverses in the product case (see Theorem 2.6). For $\nu$ real the only contribution to the limit in (5.1) comes from $\Lambda_{g}^{(N)}(\nu)$—in fact from the product case term $\Lambda_{g,0}(\nu)$. Thus the same is true for the pole of the analytic continuation of $\Lambda_{g}(\nu)$ by real analyticity. ∎
@@ A.1. Preliminaries
Proof. By [16, Prop. 4.5] (cf. [5, Thm. 6.2]), ${\mathcal{I}}_{\nu}=g_{\nu}(D)$, where $D=\frac{1}{2}(xD_{x}+D_{x}x)=i^{-1}(x\partial_{x}+\frac{1}{2})$ and $g_{\nu}$ is the real function
@@
in particular ${\rm ad}^{j}_{D}f$ is bounded if the RHS is finite. Supposing it is, we can apply the commutator expansion [15, Lem. C.3.1] (strictly speaking its direct analogue where the order of operators is reversed, which follows from the same arguments) with $\rho=0$, $N=0$ and then $N=1$, which gives
@@ A.2. Energy estimates
where $\nu(0)=\nu-1$ and $\nu(j)=\nu$ for $j=1,\dots,d$. Since $W$ is strictly time-like in the region of interest, the stress-energy tensor term involving $T^{ij}$ is negative definite and controls a homogeneous version of the $H_{\nu}^{1}$ norm of $(-\phi^{\prime})^{1/2}u$. The arguments in [21] leading to a twisted version of the Poincaré inequality [55, Prop. 2.5] remain valid, so we can also control $L^{2}$ norms and deduce the estimate
@@ Acknowledgments
The authors thank Yuchao Yi for helpful remarks. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program through the Consolidator Grant agreement 862342 (A.E.). A.E.’s research is also partially supported by the grants CEX2023-001347-S and PID2022-136795NB-I00. The research of G.U. is partly supported by NSF and a Robert R. Phelps and Elaine F. Phelps Professorship at University of Washington. M.W. gratefully acknowledges support from the grant ANR-20-CE40-0018.